# Homogenization framework for rigid and non-rigid foldable origami metamaterials

### v1

@@ 2.3.1 Theory
where $u_{\alpha}^{(\textup{r})}$, $u_{\alpha}^{(\textup{l})}$, $u_{\alpha}^{(\textup{t})}$, and $u_{\alpha}^{(\textup{b})}$ are displacements respectively on the right, left, top and bottom boundaries of the unit cell, $e_{\alpha\beta}$ are applied strains, and $L_{\beta}$ are the length of the unit cell in $y_{\beta}$ directions. The periodic boundary conditions above differ from those of asymptotic homogenization (Equation 30) by a term $e_{\alpha\beta}L_{\beta}$, which generates an average strain of $e_{\alpha\beta}$.
@@ 2.3.2 Numerical implementation
Figure 9: Step 1 of energy-based homogenization identifying deformation modes. (a) Four boundaries of the unit cell. (b) Deformed boundaries after applying periodic boundary conditions. The overall strain is $e_{11}=1$. (c) Six of the deformation modes under periodic boundary conditions with lighting effects in 3D rendering. Displacements are scaled.
@@
where $L_{1}$ and $L_{2}$ are the lengths of the unit cell in $y_{1}$ and $y_{2}$ directions, respectively.
@@ 3 Validation
where $\sigma_{\alpha\beta}$ are stresses, and $\mu_{\alpha\beta}$ are couple-stresses. The relationship involves only nine independent effective elastic constants, which are in-plane Poisson’s ratio $\nu_{12}$, Young’s moduli $E_{1}$ and $E_{2}$, shear modulus $G_{12}$, out-of-plane Poisson’s ratio $\nu^{b}_{12}$, bending moduli $M_{1}$ and $M_{2}$, and twisting modulus $T_{12}$. These relations, namely the effective elastic constants, can be determined from the effective stiffness coefficients as
@@
To set a baseline for comparison, we solved a detailed FE model of the origami sheet with identical parameters of the homogenized models, as listed in Table 1. To determine the elastic constants, we use a specimen with seven unit cells on each side of the sheet, adding up to 49 unit cells. On the other hand, for $G_{12}$ we examine a rectangular domain of 3 by 21 unit cells to avoid in-plane bending (Vasudevan and Pratapa, 2024). Although homogenization assumes an infinite tessellation of unit cells, a convergence test on the detailed FE model shows that 7 by 7 unit cells (or 3 by 21 for shear modulus) is a sufficient tessellation size. Simulations of six mechanical tests, shown in Figure 11, are required to determine all the elastic constants.
@@
Section B reports and discusses the errors with respect to the baseline results of the detailed FE models. Comparing the deviations of the plate FE asymptotic and energy-based homogenization, the bar-and-hinge model, and the rigid-panel model, we find that on average the errors are lowest for $\nu_{12}$, with the asymptotic plate FE 3.47%, the energy-based plate FE 3.20%, the bar-and-hinge 3.83%, and the rigid-panel 32.1%. The results of $G_{12}$ and $nu^{b}_{12}$ are similarly accurate. For $E_{1}$, $E_{2}$, $M_{1}$, and $M_{2}$, the averaged errors of the homogenized plate FE models are one third to one fifth those of other methods. In addition, the results for $T_{12}$ show that our framework is far more accurate (error of asymptotic plate FE 11.1%, energy-based plate FE 12.9%,) than the bar-and-hinge model (error 161%).
4 Results
We can now apply both homogenization methods to determine the effective in-plane and out-of-plane properties of Miura origami sheets with geometric and material parameters specified in Table 1. Section C provides the results in matrix form. Properties in the effective compliance matrix resemble those of a homogeneous thin plate made of an orthotropic material, with minor coupling between in-plane and out-of-plane responses. This observation confirms that the Miura pattern is centrosymmetric (Vasudevan and Pratapa, 2024).
To demonstrate the usage of the effective compliance matrices, we solve the global deformation of a Miura tessellation with seven unit cells on each side. Figure 12 shows a homogenized plate $\Omega$ with equal dimensions as the tessellation. The above compliance matrices are applied to the homogenized models as detailed in the last steps of Section 2.2.2 and Section 2.3.2. There is a fixed boundary condition on the left boundary and a load of 25 N evenly distributed on the right boundary, pointing $45^{\circ}$ to the upper right. Figures 12 and 12 presents the resulting displacement fields in the homogenized models. Identical load and boundary condition are applied to the detailed plate FE model as the baseline, whose displacement is shown in Figure 12. Both homogenized models agree well with the baseline in terms of the distribution of global displacement. There is discrepancy on the right boundary where the load is applied. Detailed model shows localized stretching of the panels that is not observed in the homogenized models.
@@ 4 Results
Figure 12: Comparison between homogenized and detailed displacement magnitudes. (a) Boundary condition and load chosen for demonstration. (b) Deformed asymptotic homogenization model.(c) Deformed energy-based homogenization model. Both homogenized models (b) and (c) have 575 elements. (d) Detailed model with 317715 elements.
4.1 Role of initial fold angle
Since the architecture of a metamaterial governs its properties, we now examine how geometry tuning impacts the effective elastic constants of a Miura sheet. In this section, the selected geometric parameter is the initial fold angle $\theta_{0}$, as shown in Figure 13. It represents the state of the pre-folded origami sheet, defined as the fold angle when the origami sheet is stress-free. We investigate the linear properties since the analysis does not consider large displacement or large strain. Among the geometric parameters listed in Table 1, we select $\theta_{0}$ as the varying parameter due to its strong influence on the folding kinematics. It also allows direct comparison with results in the literature. In this parametric study, $\theta_{0}$ varies from $1^{\circ}$ to $80^{\circ}$. Results from asymptotic and energy-based homogenization models are compared with those of the fully detailed FE model in Section 3 to determine result accuracy. After applying boundary conditions as shown in Figure 11, we characterize the detailed FE model in terms of its in-plane Poisson’s ratio $\nu_{12},$ Young’s moduli $E_{1}$, $E_{2},$ shear modulus $G_{12},$ out-of-plane Poisson’s ratio $\nu^{b}_{12},$ bending moduli $M_{1}$, $M_{2},$ and twisting modulus $T_{12}$.
Figure 13: Two configurations of Miura unit cell with distinct fold angles $\theta_{0}$.
Figures 14 and 14 compare the in-plane and out-of-plane kinematics of Miura origami through its Poisson’s ratio. The in-plane Poisson’s ratio is negative, a result that parallels the well-known observation of the auxetic response of the Miura sheet. A minimum occurs at $\theta_{0}=12^{\circ}$ with $\nu_{12}=-2.76$, while $\nu_{12}$ approaches zero when fully folded. The bending Poisson’s ratio is almost equal and opposite to the in-plane one, echoing the prediction of a pin-jointed truss framework (Schenk and Guest, 2011). These synchronized properties confirm that the Miura sheet behaves similarly to a single-DOF mechanism.
@@ 4.1 Role of initial fold angle
Figure 14: In-plane and out-of-plane effective elastic constants of Miura origami metamaterial. Effective Young’s moduli $E_{1}$, $E_{2}$, shear modulus $G_{12}$, bending moduli $M_{1}$, $M_{2}$, and twisting modulus $T_{12}$ are normalized against the base material Young’s modulus $E_{\text{base}}$. Insets show corresponding deformation modes of a detailed FE model. A subset of data points is shown for visual clarity. ( ) Asymptotic plate FE; ( ) Energy-based plate FE; (   ) Detailed plate FE.
@@
The in-plane moduli $E_{1}$, $E_{2}$, and $G_{12}$ share similar features at low to medium fold angles. For $\theta_{0}\approx 0$, the Miura sheet is almost fully flat, with properties reaching towards those of its base material, a thin sheet of Mylar. As the origami sheet folds, all effective moduli continuously decrease, because the relatively more flexible creases begin to play an increasingly influential role in the overall response. On the other hand, the kinematics of Miura pattern indicates that the total stiffness of the creases decreases for an increase in $\theta_{0}$ (Wei et al., 2013), a result that explains the trend observed on $E_{1},E_{2}$ in Figures 14 and 14. Figure 14 shows that the shear modulus $G_{12}$ is notably higher than the other two moduli. Previous work (Vasudevan and Pratapa, 2024) demonstrated that the shearing deformation mode involves mostly panel stretching, requiring a significant amount of strain energy. Finally, for $\theta_{0}>60^{\circ}$ there is a slight increase in $E_{2}$. This is explained by the alignment of panels along the $x_{2}$-direction at the highly folded state. In this case, loading along $x_{2}$ causes stretching in the panels and therefore an increase of $E_{2}$.
@@
The out-of-plane moduli $M_{1}$, $M_{2}$, and $T_{12}$ follow a similar decreasing trend for $5^{\circ}<\theta_{0}<60^{\circ}$. However, at a small $\theta_{0}$ the bending modulus $M_{1}$ in Figure 14 shows a slight increase. This effect is similar to that of a corrugated board: the flutes makes it difficult to bend in one direction but easy in the other. The Miura pattern makes z-shaped flutes along $x_{1}$ that causes an increase in $M_{1}$, but as $\theta_{0}$ increases, this effect diminishes quickly since the folding kinematics has a strong influence on $M_{1}$. The ’flutes’ in the Miura sheet also makes $M_{1}$ higher than $M_{2}$ for most $\theta_{0}$. $M_{2}$ and $T_{12}$ show a decreasing to increasing trend as $\theta_{0}$ increases, as seen in Figures 14 and 14, which is qualitatively in agreement with the response of a truss-based lattice (Vasudevan and Pratapa, 2024).
@@ 4.2 Comparison with the literature
Figure 15: Comparison of modeling methods applied to a Miura unit cell for demonstrative purposes.
@@
Figure 16: Comparison of the homogenization framework proposed in this work with methods in the literature. Sub-plots show the variation of effective elastic constants over the fold angles $\theta_{0}$ between $0^{\circ}$ and $80^{\circ}$. Except for the Poisson’s ratios $\nu_{12}$ and $\nu^{b}_{12}$, the effective elastic constants are normalized against the base material Young’s modulus $E_{\text{base}}$. Insets show zoomed-in regions of the plots. A subset of data points is shown for visual clarity. ( ) Asymptotic plate FE; ( ) Energy-based plate FE; (   ) Detailed plate FE; ( ) Homogenized/detailed bar-and-hinge; ( ) Rigid panel.
For $0^{\circ}<\theta<10^{\circ}$, the trend of both numerical models (bar-and-hinge and plate FE) differ from that of the rigid-panel model, with the former decreasing and the latter increasing. The monotonic increase in the rigid-panel model originates from the assumption that the origami metamaterial behaves as a 1-DoF mechanism. Its kinematic relation fully prescribes the mechanical response in Equation E.1. On the other hand, the two numerical models account for the base material properties. When $\theta$ approaches 0, the origami metamaterial unfolds to a flat sheet of the base material, i.e., Mylar in this work. Correspondingly, the Poisson’s ratio from numerical models approaches that of Mylar. As $\theta$ varies, the base material properties and origami kinematics play a competing role in the metamaterial response. With the increase of $\theta$, the kinematic relationship gradually takes dominance. The Poisson’s ratio first decreases and then increases to align with the response of a pure mechanism, forming a local minimum at $\theta=10^{\circ}$ under the given parameters. Additionally, both the plate FE model and the bar-and-hinge model confirm that $\nu^{b}_{12}$ in Figure 16 is close to the negative of $\nu_{12}$ in Figure 16.
Figures 16 and 16 show that the in-plane moduli of all methods share the same trend. The Young’s moduli $E_{1},E_{2}$ of the plate FE model are lower than others. For example, at a given fold angle $\theta_{0}=30^{\circ}$, the moduli of the bar-and-hinge model are close to 1.5 times those of the plate FE. The discrepancy grows as $\theta_{0}$ becomes smaller and the Miura origami unfolds towards a flat sheet. The bar-and-hinge and the rigid-panel models align well with each other for most of the cases, when $10^{\circ}<\theta_{0}<80^{\circ}$. The shear modulus $G_{12}$ of plate FE and bar-and hinge models match in their values. The rigid-panel model is not included since it does not allow shearing.

### v3

@@ 2.3.1 Theory
where $u_{\alpha}^{(\textup{r})}$, $u_{\alpha}^{(\textup{l})}$, $u_{\alpha}^{(\textup{t})}$, and $u_{\alpha}^{(\textup{b})}$ are displacements respectively on the right, left, top and bottom boundaries of the unit cell, $e_{\alpha\beta}$ are applied strains, and $L_{\beta}$ are the length of the unit cell along $y_{\beta}$. The periodic boundary conditions above differ from those of asymptotic homogenization (Equation 30) by a term $e_{\alpha\beta}L_{\beta}$, which generates an average strain of $e_{\alpha\beta}$.
@@ 2.3.2 Numerical implementation
Figure 9: Step 1 of energy-based homogenization identifying deformation modes. (a) Four boundaries of the unit cell. (b) Deformed boundaries after applying periodic boundary conditions. The overall strain is $e_{11}=1$. (c) Six of the deformation modes under periodic boundary conditions with lighting effects in 3D rendering. Displacements are down-scaled for plotting purposes.
@@
where $L_{1}$ and $L_{2}$ are the lengths of the unit cell along $y_{1}$ and $y_{2}$, respectively.
@@ 3 Validation
where $\sigma_{\alpha\beta}$ are stresses, and $\mu_{\alpha\beta}$ are couple-stresses. The relationship involves only nine independent effective elastic constants, which are in-plane Poisson’s ratio $\nu_{12}$, Young’s moduli $E_{1}$ and $E_{2}$, shear modulus $G_{12}$, out-of-plane Poisson’s ratio $\nu^{b}_{12}$, bending moduli $M_{1}$ and $M_{2}$, and twisting modulus $T_{12}$. The out-of-plane Poisson’s ratio $\nu^{b}_{12}$ and $\nu^{b}_{21}$ are defined from the transverse curvatures, $\nu^{b}_{12}=-\kappa_{22}/\kappa_{11}$, and $\nu^{b}_{21}=-\kappa_{11}/\kappa_{22}$.
The effective elastic constants can be determined from the effective stiffness coefficients as
@@
To set a baseline for comparison, we solved a detailed FE model of the origami sheet with identical parameters of the homogenized models, as listed in Table 1. Standard practice in homogenization applied to both periodic porous composites (Hollister and Kikuchi, 1992) and origami metamaterials (Vasudevan and Pratapa, 2024) shows that a detailed FE model is sufficient to validate the result from homogenization. To determine the elastic constants, we use a specimen with seven unit cells on each side of the sheet, adding up to 49 unit cells. On the other hand, for $G_{12}$ we examine a rectangular domain of 3 by 21 unit cells to avoid in-plane bending (Vasudevan and Pratapa, 2024). Although homogenization assumes an infinite tessellation of unit cells, a convergence test on the detailed FE model shows that 7 by 7 unit cells (or 3 by 21 for shear modulus) is a sufficient tessellation size. Simulations of six mechanical tests, shown in Figure 11, are required to determine all the elastic constants. To reduce the boundary effect, we measure the strain or curvature of the unit cell at the center of the tessellation and compute the reference elastic constants.
@@
[…]
Figure 12: Experimental, homogenized and detailed simulation of force-displacement curve of Miura origami. (a) Homogenized model under roller boundary condition and compression load. (b) Detailed model of Miura origami. (c) Specimen used for experiment. (d) Compression test setup. (e) Force-displacement curve comparing experiment, homogenized model and detailed model. The experimental curve consists of a shade indicating the range of 6 tests, as well as a solid curve showing a representative test.
4 Results
We can now apply both homogenization methods to determine the effective in-plane and out-of-plane properties of Miura origami sheets. Properties in the effective compliance matrix resemble those of a homogeneous thin plate made of an orthotropic material, with minor coupling between in-plane and out-of-plane responses. This observation confirms that the Miura pattern is centrosymmetric (Vasudevan and Pratapa, 2024).
@@ 4 Results
[…]
Figure 13: Comparison of homogenized displacement fields against detailed simulation and experiment. (a) Boundary condition and load chosen for demonstration on homogenized model. (b) $u_{1}$ displacements of asymptotic homogenization model.(c) $u_{2}$ displacements of asymptotic homogenization model. (d) $u_{1}$ displacements of energy-based homogenization model. (e) $u_{2}$ displacements of energy-based homogenization model. The homogenized models in (b)-(e) have 575 elements. (f) Equivalent load and boundary condition on detailed model with 317,715 elements. (g) $u_{1}$ displacements of detailed model. (h) $u_{2}$ displacements of detailed model. (i) Experimental setup. A cardboard Miura origami with left edge glued to a fixture and right edge attached to a slider. (j) Deformed specimen under applied load on the slider. All simulation results are superimposed on a grid with a cell size equal to that used in the experiment.
[…]
To demonstrate the applicability range of the framework across fold patterns, we report in Section G the effective properties of a curve crease origami pattern adapted from the literature (Liu and James, 2024).
4.1 Role of initial fold angle
Since the architecture of a metamaterial governs its properties, we now examine how geometry tuning impacts the effective elastic constants of a Miura sheet. In this section, the selected geometric parameter is the initial fold angle $\theta_{0}$, as shown in Figure 14. It is measured as the dihedral angle between the panel and the $x_{1}$-$x_{2}$ plane. All four panels have the identical angle, a result stemming from the kinematics of Miura origami. $\theta_{0}$ represents the state of the pre-folded origami sheet, defined as the fold angle when the origami sheet is stress-free. We investigate the linear properties since the analysis does not consider large displacement or large strain. Among the geometric parameters listed in Table 1, we select $\theta_{0}$ as the varying parameter due to its strong influence on the folding kinematics. It also allows direct comparison with the bar-and-hinge model (Vasudevan and Pratapa, 2024) and the rigid-panel model (Wei et al., 2013), both of which have computed the effective properties of the Miura pattern under various fold angles. In this parametric study, $\theta_{0}$ varies from $1^{\circ}$ to $80^{\circ}$. Results from asymptotic and energy-based homogenization models are compared with those of the fully detailed FE model in Section 3 to determine result accuracy. After applying boundary conditions as shown in Figure 11, we characterize the detailed FE model in terms of its in-plane Poisson’s ratio $\nu_{12},$ Young’s moduli $E_{1}$, $E_{2},$ shear modulus $G_{12},$ out-of-plane Poisson’s ratio $\nu^{b}_{12},$ bending moduli $M_{1}$, $M_{2},$ and twisting modulus $T_{12}$. To reduce the boundary effect, we measure the above elastic constants using the unit cell at the center of the origami metamaterial.
Figure 14: Three configurations of Miura unit cell with increasing fold angle $\theta_{0}$.
Figures 15 and 15 compare the in-plane and out-of-plane kinematics of Miura origami through its Poisson’s ratio. The in-plane Poisson’s ratio is negative, a result that parallels the well-known observation of the auxetic response of the Miura sheet. A minimum occurs at $\theta_{0}=12^{\circ}$ with $\nu_{12}=-2.76$, while $\nu_{12}$ approaches zero when fully folded. The bending Poisson’s ratio is almost equal and opposite to the in-plane one, echoing the prediction of a pin-jointed truss framework (Schenk and Guest, 2011). These synchronized properties confirm that the Miura sheet behaves similarly to a single-DOF mechanism.
[…]
@@ 4.1 Role of initial fold angle
Figure 15: In-plane and out-of-plane effective elastic constants of Miura origami metamaterial, including (a) in-plane Poisson’s ratio $\nu_{12}$, effective Young’s moduli (b) $E_{1}$, (c) $E_{2}$, (d) shear modulus $G_{12}$, (e) out-of-plane Poisson’s ratio $\nu^{b}_{12}$, bending moduli (f) $M_{1}$, (g) $M_{2}$, and (h) twisting modulus $T_{12}$. $E_{1}$, $E_{2}$, $G_{12}$, $M_{1}$, $M_{2}$, and $T_{12}$ are normalized against the base material Young’s modulus $E_{\text{base}}$. Insets show corresponding deformation modes of a detailed FE model. A subset of data points is shown for visual clarity. ( ) Asymptotic plate FE; ( ) Energy-based plate FE; (   ) Detailed plate FE.
@@
The in-plane moduli $E_{1}$, $E_{2}$, and $G_{12}$ share similar features at low to medium fold angles. For $\theta_{0}\approx 0^{\circ}$, the Miura sheet is almost fully flat, with properties reaching towards those of its base material, a thin sheet of Mylar. As the origami sheet folds, all effective moduli continuously decrease, because the relatively more flexible creases begin to play an increasingly influential role in the overall response. On the other hand, the kinematics of Miura pattern indicates that the total stiffness of the creases decreases for an increase in $\theta_{0}$ (Wei et al., 2013), a result that explains the trend observed on $E_{1},E_{2}$ in Figures 15 and 15. Figure 15 shows that the shear modulus $G_{12}$ is notably higher than the other two moduli. Previous work (Vasudevan and Pratapa, 2024) demonstrated that the shearing deformation mode involves mostly panel stretching, requiring a significant amount of strain energy. Finally, for $\theta_{0}>60^{\circ}$ there is a slight increase in $E_{2}$. This is explained by the alignment of panels along the $x_{2}$-direction at the highly folded state. In this case, loading along $x_{2}$ causes stretching in the panels and therefore an increase of $E_{2}$.
@@
The out-of-plane moduli $M_{1}$, $M_{2}$, and $T_{12}$ follow a similar decreasing trend for $5^{\circ}<\theta_{0}<60^{\circ}$. However, at a small $\theta_{0}$ the bending modulus $M_{1}$ in Figure 15 shows a slight increase. This effect is similar to that of a corrugated board: the flutes makes it difficult to bend in one direction but easy in the other. The Miura pattern makes z-shaped flutes along $x_{1}$ that causes an increase in $M_{1}$, but as $\theta_{0}$ increases, this effect diminishes quickly since the folding kinematics has a strong influence on $M_{1}$. The ’flutes’ in the Miura sheet also makes $M_{1}$ higher than $M_{2}$ for most $\theta_{0}$. $M_{2}$ and $T_{12}$ show a decreasing to increasing trend as $\theta_{0}$ increases, as seen in Figures 15 and 15, which is qualitatively in agreement with the response of a truss-based lattice (Vasudevan and Pratapa, 2024).
@@ 4.2 Comparison with the literature
Figure 16: Comparison of modeling methods applied to a Miura unit cell for demonstrative purposes. (a) Rigid panel model has ideally rigid panels and creases represented with torsional springs. (b) Bar-and-hinge model has crease bars to model crease folding and diagonal bars to model panel bending. (c) Plate FE model has torsional springs as creases and panels meshed with plate elements.
@@
Figure 17: Comparison of the homogenization framework proposed in this work with methods in the literature. Sub-plots show the variation of effective elastic constants over the fold angles $\theta_{0}$ between $0^{\circ}$ and $80^{\circ}$, including (a) in-plane Poisson’s ratio $\nu_{12}$, effective Young’s moduli (b) $E_{1}$, (c) $E_{2}$, (d) shear modulus $G_{12}$, (e) out-of-plane Poisson’s ratio $\nu^{b}_{12}$, bending moduli (f) $M_{1}$, (g) $M_{2}$, and (h) twisting modulus $T_{12}$. $E_{1}$, $E_{2}$, $G_{12}$, $M_{1}$, $M_{2}$, and $T_{12}$ are normalized against the base material Young’s modulus $E_{\text{base}}$. Insets show zoomed-in regions of the plots. A subset of data points is shown for visual clarity. ( ) Asymptotic plate FE; ( ) Energy-based plate FE; (   ) Detailed plate FE; ( ) Homogenized/detailed bar-and-hinge; ( ) Rigid panel.
For $0^{\circ}<\theta<10^{\circ}$, the trend of both numerical models (bar-and-hinge and plate FE) differ from that of the rigid-panel model, with the former decreasing and the latter increasing. The monotonic increase in the rigid-panel model originates from the assumption that the origami metamaterial behaves as a 1-DoF mechanism. Its kinematic relation fully prescribes the mechanical response in Equation E.1. On the other hand, the two numerical models account for the base material properties. When $\theta$ approaches $0^{\circ}$, the origami metamaterial unfolds to a flat sheet of the base material, i.e., Mylar in this work. Correspondingly, the Poisson’s ratio from numerical models approaches that of Mylar. As $\theta$ varies, the base material properties and origami kinematics play a competing role in the metamaterial response. With the increase of $\theta$, the kinematic relationship gradually takes dominance. The Poisson’s ratio first decreases and then increases to align with the response of a pure mechanism, forming a local minimum at $\theta=10^{\circ}$ under the given parameters. Additionally, both the plate FE model and the bar-and-hinge model confirm that $\nu^{b}_{12}$ in Figure 17 is close to the negative of $\nu_{12}$ in Figure 17.
Figures 17 and 17 show that the in-plane moduli of all methods share the same trend. The Young’s moduli $E_{1},E_{2}$ of the plate FE model are lower than others. For example, at a given fold angle $\theta_{0}=30^{\circ}$, the moduli of the bar-and-hinge model are close to 1.5 times those of the plate FE. The discrepancy grows as $\theta_{0}$ becomes smaller and the Miura origami unfolds towards a flat sheet. The bar-and-hinge and the rigid-panel models align well with each other for most of the cases, when $10^{\circ}<\theta_{0}<80^{\circ}$. The shear modulus $G_{12}$ of plate FE and bar-and hinge models match in their values. The rigid-panel model is not included since it does not allow shearing.